### v1

@@ 7 Discussion
In this work, we introduced the notion of short certificates: subsets of the training data that provably determine the correct label of a test point $x$ even under up to $b$ corruptions. To characterize their worst-case size, we proposed the robust hollow star number, a generalization of the hollow star number of [5]. We then studied worst-case distributional bounds and introduced the certificate coefficient $\varepsilon_{x}$, which captures the distribution-dependent difficulty of certifying a given point. For this setting, we established bounds on the sample size required to certify $x$, as a function of $\varepsilon_{x}$, the corruption budget $b$, and the VC-dimension $d$ of the class $\mathcal{H}$. We further showed that reweighted variants of $\varepsilon_{x}$, can lead to improved bounds from polynomial-sized samples. Our framework also naturally subsumes the agnostic learning setting as a special case of adversarial model, under the assumption that the true label of $x$ is given by the hypothesis closest to the target function. Following [2], a $b$-robust certificate becomes a $b$-agnostic certificate when the corruption budget reflects the error rate of the best-in-class hypothesis: a set $S$ certifies $(x,y)$ if every hypothesis predicting $1-y$ on $x$ incurs more than $b$ errors on $S$. We believe Open Question 3.9 and Open Question 6.5 pose exciting directions for future research.
@@ Acknowledgments
This work was supported in part by the National Science Foundation under grants CCF-2212968, ECCS-2216899, and ECCS-2217023, by the Simons Foundation under the Simons Collaboration on the Theory of Algorithmic Fairness, and by the Office of Naval Research MURI Grant N000142412742.
@@ A.2 $b$-robust Hollow Star Number for Singletons
Consider the class $\mathcal{H}$ of singletons on a domain of size $n$, i.e., $\mathcal{X}=[n]$, =$\mathcal{H}=\{x\mapsto(-1)^{1-\mathds{1}[x=i]}:i\in[n]\}$. The $b$-robust hollow star number of $\mathcal{H}$ is exactly equal to $(b+1)(n-1)+1$.
@@
Suppose that $T=(x_{1},y_{1},b+1),(x_{2},y_{2},1),\dots,(x_{s_{b}},y_{s_{b}},1)$ is a $b$-robust hollow star for $\mathcal{H}$, where as per Definition 3.4, the weight on $(x_{1},y_{1})$ is $b+1$, and the weight on the rest of the points is $1$. Note that no matter what $y_{1}$ is, there exists some $h_{i^{\prime}}\in\mathcal{H}$ that labels $x_{1}$ as $-y_{1}$.
Now, consider the sequence $T\setminus\{(x_{s_{b}},y_{s_{b}},1)\}$. There should exist $h_{i_{1}}\in\mathcal{H}$ that $b$-robustly realizes this sequence. In particular, $h_{i_{1}}$ must label $x_{1}$ as $y_{1}$ (hence, $h_{i_{1}}$ cannot be $h_{i^{\prime}}$). Furthermore, it must be the case that $h_{i_{1}}$ labels $x_{s_{b}}$ as $-y_{s_{b}}$; otherwise, $h_{i_{1}}$ would $b$-robustly realize all of $T$. Thus, $h_{i_{1}}$ is a hypothesis in $\mathcal{H}$ which: (1) is not $h_{i^{\prime}}$, (2) makes a mistake on $(x_{s_{b}},y_{s_{b}})$, and (3) makes at most $b$ mistakes on $(x_{2},y_{2}),\dots,(x_{s_{b}-1},y_{s_{b}-1})$.
@@
Thus, we can think of the $n$ hypotheses in $\mathcal{H}$ as $n$ bins, and the $s_{b}-1$ points $(x_{2},y_{2}),\dots,(x_{s_{b}},y_{s_{b}})$ as $s_{b}-1$ balls. From conditions (1) and (2), we get that each of the $s_{b}-1$ balls is assigned one of $n-1$ bins from the set $\mathcal{H}\setminus\{h_{i^{\prime}}\}$. From condition (3), each of the $n-1$ bins is assigned no more than $b+1$ balls. Thus, it must hold that $(b+1)(n-1)\geq s_{b}-1$, which completes the proof.
@@
$=\operatorname*{\mathbb{P}}_{S^{\prime\prime}}[\exists h\in\mathcal{H}_{x}|_{S^{\prime\prime}}\text{ s.t. $h$ makes at most $\varepsilon_{x}m/4$ mistakes on $S$, but at least $\varepsilon_{x}m/2$ mistakes on $S^{\prime}$}]$
$\leq\sum_{h\in\mathcal{H}_{x}|_{S^{\prime\prime}}}\operatorname*{\mathbb{P}}_{S^{\prime\prime}}[\text{$h$ makes at most $\varepsilon_{x}m/4$ mistakes on $S$, but at least $\varepsilon_{x}m/2$ mistakes on $S^{\prime}$}].$
Fix any $h\in\mathcal{H}_{x}|_{S^{\prime\prime}}$. In the form above, once $S^{\prime\prime}$ is drawn, the event in the parentheses either happens or it doesn’t. So, we introduce some additional randomness. Suppose
@@
We now flip $m$ independent fair coins $p_{1},\dots,p_{m}$. If $p_{i}$ lands heads, we swap $(x_{i},h^{\star}(x_{i}))$ and $(x_{m+i},h^{\star}(x_{m+i}))$, otherwise we let this pair be as is.
After doing this for all the $m$ coins, we let the first $m$ points be $S$ and the second $m$ points be $S^{\prime}$. Again, this is a distributionally identical way of obtaining $S$ and $S^{\prime}$. So, want to bound
$\operatorname*{\mathbb{P}}_{S^{\prime\prime}}\operatorname*{\mathbb{P}}_{p_{1},\dots,p_{m}}[\text{$h$ makes at most $\varepsilon_{x}m/4$ mistakes on $S$, but at least $\varepsilon_{x}m/2$ mistakes on $S^{\prime}$}].$
Now, after drawing $S^{\prime\prime}$ (and conditioning on it/fixing it), the event in the parentheses depends on how the coins $p_{1},\ldots,p_{m}$ pan out. So, for any fixed $S^{\prime\prime}$, we will bound the probability of the event in the parentheses, where the randomness is only over the coin flips.
@@
$\operatorname*{\mathbb{P}}[A]\leq 2\operatorname*{\mathbb{P}}[B]\leq 2\sum_{h\in\mathcal{H}_{x}|_{S^{\prime\prime}}}\delta/2\tau_{\mathcal{H}}(2m)\leq\tau_{\mathcal{H}}(2m)\cdot\delta/2\tau_{\mathcal{H}}(2m)\leq\delta,$
@@
We will show that drawing $\mbox{\rm poly}(b,d,1/\varepsilon_{x},1/\delta)$ samples from $\mathcal{D}$ is sufficient to ensure that with probability $21-\delta/$ 2, the number of samples accepted is at least $6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)$. Conditioned on this event, note that all the accepted samples are distributed according to $\mathcal{D}_{w}$. Denote the first $m_{w}=6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)$ of the accepted samples as $S^{\prime}$. Then, Theorem 5.2 immediately gives us that with probability at least $21-\delta/$ 2, $(x,h^{\star}(x))$ belongs to the $b$-robust agreement region of $S^{\prime}$. A union bound over both the $\delta/2$ failure probability events ensures that $S^{\prime}$ thus obtained is the required certificate from the theorem statement, with probability $1-\delta$. The size of $S^{\prime}$ is $m_{w}=6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)\leq 12\left(\frac{b+d\log(2/\varepsilon^{\star}_{x})+\log(2/\delta)}{\varepsilon^{\star}_{x}}\right)$ as required.
[…]
@@
where in the first inequality, we used the condition that $w$ is a valid reweighting (4). Therefore, for this setting of $m$, with probability at least $21-\delta/$ 2, the number of accepted samples is at least

### v4

@@ 7 Discussion
In this work, we introduced the notion of short certificates: subsets of the training data that provably determine the correct label of a test point $x$ even under up to $b$ corruptions. To characterize their worst-case size, we proposed the robust hollow star number, a generalization of the hollow star number of BHMZ (20). We then studied worst-case distributional bounds and introduced the certificate coefficient $\varepsilon_{x}$, which captures the distribution-dependent difficulty of certifying a given point. For this setting, we established bounds on the sample size required to certify $x$, as a function of $\varepsilon_{x}$, the corruption budget $b$, and the VC-dimension $d$ of the class $\mathcal{H}$. We further showed that reweighted variants of $\varepsilon_{x}$, can lead to improved bounds from polynomial-sized samples. Our framework also naturally subsumes the agnostic learning setting as a special case of adversarial model, under the assumption that the true label of $x$ is given by the hypothesis closest to the target function. Following BBHS (22), a $b$-robust certificate becomes a $b$-agnostic certificate when the corruption budget reflects the error rate of the best-in-class hypothesis: a set $S$ certifies $(x,y)$ if every hypothesis predicting 1y1-y on $x$ incurs more than $b$ errors on $S$. We believe Open Question 6.5 poses an exciting direction for future research.
@@ Acknowledgments
This work was supported in part by the National Science Foundation under grants CCF-2212968, ECCS-2216899, and ECCS-2217023, by the Simons Foundation under the Simons Collaboration on the Theory of Algorithmic Fairness, and by the Office of Naval Research MURI Grant N000142412742. We thank Shay Moran for bringing to our notice the work by Bárány and Perles BP (90), which resolves Open Question 3.9.
@@ A.2 $b$-robust Hollow Star Number for Singletons
Consider the class $\mathcal{H}$ of singletons on a domain of size $n$, i.e., $\mathcal{X}=[n]$, $\mathcal{H}=\{x\mapsto(-1)^{1-\mathds{1}[x=i]}:i\in[n]\}$. The $b$-robust hollow star number of $\mathcal{H}$ is exactly equal to (b+1)(n1)+1(b+1)(n-1)+1.
@@
Suppose that $T=(x_{1},y_{1},b+1),(x_{2},y_{2},1),\dots,(x_{s_{b}},y_{s_{b}},1)$ is a $b$-robust hollow star for $\mathcal{H}$, where as per Definition 3.4, the weight on $(x_{1},y_{1})$ is $b+1$, and the weight on the rest of the points is $1$. Note that no matter what $y_{1}$ is, there exists some $h_{i^{\prime}}\in\mathcal{H}$ that labels $x_{1}$ as $y_{1}$.
Now, consider the sequence $T\setminus\{(x_{s_{b}},y_{s_{b}},1)\}$. There should exist $h_{i_{1}}\in\mathcal{H}$ that $b$-robustly realizes this sequence. In particular, $h_{i_{1}}$ must label $x_{1}$ as $y_{1}$ (hence, $h_{i_{1}}$ cannot be $h_{i^{\prime}}$). Furthermore, it must be the case that $h_{i_{1}}$ labels $x_{s_{b}}$ as $y_{s_{b}}$; otherwise, $h_{i_{1}}$ would $b$-robustly realize all of $T$. Thus, $h_{i_{1}}$ is a hypothesis in $\mathcal{H}$ which: (1) is not $h_{i^{\prime}}$, (2) makes a mistake on $(x_{s_{b}},y_{s_{b}})$, and (3) makes at most $b$ mistakes on $(x_{2},y_{2}),\dots,(x_{s_{b}-1},y_{s_{b}-1})$.
@@
Thus, we can think of the $n$ hypotheses in $\mathcal{H}$ as $n$ bins, and the $s_{b}-1$ points $(x_{2},y_{2}),\dots,(x_{s_{b}},y_{s_{b}})$ as $s_{b}-1$ balls. From conditions (1) and (2), we get that each of the $s_{b}-1$ balls is assigned one of n1n-1 bins from the set $\mathcal{H}\setminus\{h_{i^{\prime}}\}$. From condition (3), each of the n1n-1 bins is assigned no more than $b+1$ balls. Thus, it must hold that $(b+1)(n-1)\geq s_{b}-1$, which completes the proof.
@@
[…]
In the form above, once $S^{\prime\prime}$ is drawn, the event in the parentheses is determined. So, we introduce some additional randomness. Suppose
@@
We now flip $m$ independent fair coins $p_{1},\dots,p_{m}$. If $p_{i}$ lands heads, we swap $(x_{i},h^{\star}(x_{i}))$ and $(x_{m+i},h^{\star}(x_{m+i}))$, otherwise we let this pair be as is. After doing this for all the $m$ coins, we let the first $m$ points be $S$ and the second $m$ points be $S^{\prime}$. Again, this is a distributionally identical way of obtaining $S$ and $S^{\prime}$. So, we have that
[…]
Now, for any fixed $S^{\prime\prime}$ and $h\in\mathcal{H}_{x}|_{S^{\prime\prime}}$, we will bound the probability of the event in the parentheses, where the randomness is only over the coin flips.
@@
$\operatorname*{\mathbb{P}}[A]\leq 2\operatorname*{\mathbb{P}}[B]\leq 2\operatorname*{{\mathbb{E}}}_{S^{\prime\prime}}\left[\sum_{h\in\mathcal{H}_{x}|_{S^{\prime\prime}}}\delta/2\tau_{\mathcal{H}}(2m)\right]\leq 2\tau_{\mathcal{H}}(2m)\cdot\delta/2\tau_{\mathcal{H}}(2m)=\delta,$
@@
We will show that drawing $\mbox{\rm poly}(b,d,1/\varepsilon_{x},1/\delta)$ samples from $\mathcal{D}$ is sufficient to ensure that with probability $1-\delta/2$, the number of samples accepted is at least $6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)$. Conditioned on this event, note that all the accepted samples are distributed according to $\mathcal{D}_{w}$. Denote the first $m_{w}=6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)$ of the accepted samples as $S^{\prime}$. Then, Theorem 5.2 immediately gives us that with probability at least $1-\delta/2$, $(x,h^{\star}(x))$ belongs to the $b$-robust agreement region of $S^{\prime}$. A union bound over both the $\delta/2$ failure probability events ensures that $S^{\prime}$ thus obtained is the required certificate from the theorem statement, with probability $1-\delta$. The size of $S^{\prime}$ is $m_{w}=6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)\leq 12\left(\frac{b+d\log(2/\varepsilon^{\star}_{x})+\log(2/\delta)}{\varepsilon^{\star}_{x}}\right)$ as required.
So, we continue to argue that if we draw $\mbox{\rm poly}(b,d,1/\varepsilon_{x},1/\delta)$ samples from $\mathcal{D}$, then at least $6\left(\frac{b+d\log(1/\varepsilon^{w}_{x})+\log(2/\delta)}{\varepsilon^{w}_{x}}\right)$ samples are accepted. Towards this, note that by definition of the sampling process, a sample drawn from $\mathcal{D}$ gets accepted with probability $p=\int_{z}w(z)\mathcal{D}(z)dz$. Thus, if we draw $m$ samples i.i.d. from $\mathcal{D}$, the expected number of accepted samples is $pm$. By a Chernoff bound, the probability that less than $pm/2$ samples get accepted is at most $\exp(-pm/12)$, which is at most $\delta/2$ provided $m\geq\frac{12\log(2/\delta)}{p}$. Thus, for $m$ satisfying this condition, we have that with probability at least $1-\delta/2$, at least $pm/2$ samples are accepted.
@@
where in the first inequality, we used the condition that $w$ is a valid reweighting (4). Therefore, for this setting of $m$, with probability at least $1-\delta/2$, the number of accepted samples is at least